# Each to Their Own: Exploring the Optimal Embedding in RAG

### v2

@@ Abstract
Recently, as Large Language Models (LLMs) have fundamentally impacted various fields, the methods for incorporating up-to-date information into LLMs or adding external knowledge to construct domain-specific models have garnered wide attention. Retrieval-Augmented Generation (RAG), serving as an inference-time scaling method, is notable for its low cost and minimal effort for parameter tuning. However, due to heterogeneous training data and model architecture, the variant embedding models used in RAG exhibit different benefits across various areas, often leading to different similarity calculation results and, consequently, varying response quality from LLMs. To address this problem, we propose and examine two approaches to enhance RAG by combining the benefits of multiple embedding models, named Mixture-Embedding RAG and Confident RAG. Mixture-Embedding RAG simply sorts and selects retrievals from multiple embedding models based on standardized similarity; however, it does not outperform vanilla RAG. In contrast, Confident RAG generates responses multiple times using different embedding models and then selects the responses with the highest confidence level, demonstrating average improvements of approximately 10% and 5% over vanilla LLMs and RAG, respectively. The consistent results across different LLMs and embedding models indicate that Confident RAG is an efficient plug-and-play approach for various domains. We will release our code upon publication.
1 Introduction
Large language models (LLMs) have recently accelerated the pace of transformation across multiple fields, including transportation Lyu et al. (2025), arts Zhao et al. (2025), and education Gao et al. (2024), through various paradigms such as direct answer generation, training from scratch on different types of data, and fine-tuning on target domains. However, the hallucination problem Henkel et al. (2024) associated with LLMs has confused people for a long time, stemming from multiple factors such as a lack of knowledge on the given prompt Huang et al. (2025b) and a biased training process Zhao (2025).
@@ 1 Introduction
Serving as a highly efficient solution, Retrieval-Augmented Generation (RAG) has been widely employed in constructing foundation models Chen et al. (2024) and practical agents Arslan et al. (2024). Compared to training methods like fine-tuning and prompt-tuning, its plug-and-play feature makes RAG an efficient, simple, and cost-effective approach. The main paradigm of RAG involves first calculating the similarities between a question and chunks in an external knowledge corpus, followed by incorporating the top $K$ relevant chunks into the prompt to guide the LLMs Lewis et al. (2020).
Despite the advantages of RAG, selecting the appropriate embedding models remains a crucial concern, as the quality of retrieved references directly influences the generation results of the LLM Tu et al. (2025). Variations in training data and model architecture lead to different embedding models providing benefits across various domains. The differing similarity calculations across embedding models often leave researchers uncertain about how to choose the optimal one. Consequently, improving the accuracy of RAG from the perspective of embedding models continues to be an ongoing area of research.
@@
To address this research gap, we propose two methods for improving RAG by combining the benefits of multiple embedding models. The first method is named Mixture-Embedding RAG, which sorts the retrieved materials from multiple embedding models based on normalized similarity and selects the top $K$ materials as final references. The second method is named Confident RAG, where we first utilize vanilla RAG to generate answers multiple times, each time employing a different embedding model and recording the associated confidence metrics, and then select the answer with the highest confidence level as the final response. By validating our approach using multiple LLMs and embedding models, we illustrate the superior performance and generalization of Confident RAG, even though Mixture-Embedding RAG may lose to vanilla RAG. The main contributions of this paper can be summarized as follows:
@@
  • We first point out that in RAG, different embedding models operate within their own prior domains. To leverage the strengths of various embedding models, we propose and test two novel RAG methods: Mixture-Embedding RAG and Confident RAG. These methods effectively utilize the retrieved results from different embedding models to their fullest extent.
@@
2 Preliminary and Related Work
2.1 Retrieval Augmented Generation (RAG)
@@
Previous studies have validated a wide range of modifications to the RAG pipeline, demonstrating effectiveness in open-domain question answering by improving various components such as adaptive retrieval strategies Tang et al. (2024); Guan et al. (2024), dynamic document selection Kim and Lee (2024); Hei et al. (2024), joint retriever-generator optimization Siriwardhana et al. (2021), reinforcement learning Huang et al. (2025a), and hybrid retrieval-reranking frameworks Sawarkar et al. (2024); Zhang et al. (2025).
@@
Since embedding models can produce different similarity calculations, many studies have innovated in retriever embedding models to improve the performance of the RAG pipeline. For example, Invar-RAG Liu et al. (2024) constructs an LLM-based retriever with LoRA-based representation learning and an invariance loss, directly addressing retriever embedding variance and locality, and demonstrates improved retrieval and response accuracy. Blended RAG Sawarkar et al. (2024) combines dense semantic embeddings with sparse retrievers, empirically showing that this hybrid embedding space achieves higher QA accuracy, especially as corpus size scales. W-RAG Nian et al. (2024) further validates that enhancing retriever embeddings using weak signals from the downstream QA task offers gains comparable to methods requiring full human annotations.
However, most studies have focused on the innovation of a single embedding model. Due to variations in model architecture and training datasets, different embedding models operate effectively in different domains. Our approach will leverage the strengths of various embedding models to provide the most relevant information for LLMs.
2.2 Confidence of LLMs
@@
The confidence level of the LLM output can indicate the response quality to some extent. Several approaches have been employed to measure model confidence. Some of them depend on model prompting, external similarity or model-expressed judgments. These include linguistic confidence measures Shrivastava et al. (2023); Xiong et al. (2023), where the model is prompted to state its confidence explicitly as part of the output, as well as LLM-as-judge or self-evaluation strategies that involve prompting the model to rate or compare its own answers in scalar or relative terms Shrivastava et al. (2025); Ren et al. (2023); Hager et al. (2025). However, these methods are often poorly calibrated, not directly related to the true predictive uncertainty of the model, and can be biased or inconsistent due to their reliance on language generation rather than statistical likelihood.
@@
In contrast, probability-based metrics directly ground into models’ token probability output Chen and Mueller (2023); Jiang et al. (2021); Lin et al. (2024). These include average log-probability as the mean token log-probabilities, entropy-based measures that quantify prediction uncertainty via sequence or per-token entropy, self-certainty Kang et al. (2025) that utilizes the intrinsic probability distribution of LLM outputs to assess the model confidence.
@@
Many studies have observed that the probability of each token reflects the LLMs’ confidence. Intuitively, if the selected token has a higher chosen probability $p(y_{i}|x,y_{<i})$, or if the generated probability vector $\text{f}(x,y_{<i})$ exhibits a less uniform distribution, the LLMs demonstrate higher confidence in their generated results. Our study will use the confidence measurement metrics summarized in Kang et al. (2025):
@@
3 Methodology
@@
Based on the understanding that different embedding models excel in their respective domains, our study aims to develop a comprehensive RAG method that combines the strengths of various embedding models. In this work, given a question $q$, we utilize a single LLM $\text{f}()$ for answer generation, while NN embedding models $\{\text{g}_{1}(),\text{g}_{2}(),\ldots,\text{g}_{N}()\}$ assist in retrieving question-related information from a corpus, which has been divided into chunks $C=[c_{1},c_{2},\ldots,c_{m}]$. Inspired by the stacking and bagging paradigms of ensemble learning, we propose two distinct methods: (1) Mixture-Embedding RAG, which first combines the retrieved results from multiple embedding models and incorporates them into the prompt for the LLM, and (2) Confident RAG, which employs vanilla RAG with different embedding models for initial answer generation and subsequently selects the answer with the highest confidence as the final response.
@@
As shown in Fig. 1, the two proposed methods share the same preliminary process: first, they utilize different embedding models to identify the $k$ chunks with the highest similarity, similar to the approach taken by the vanilla RAG described in Section 2.1. We define the retrieved chunks as $R=\{r_{i,j}\mid i=1,2,\ldots,N;j=1,2,\ldots,k\}$, where $r_{i,j}$ is defined as $c_{\mathcal{K}_{j}^{i}}$ and $\mathcal{K}^{i}$ is the index set of the selected $k$ chunks by embedding model $\text{g}_{i}$. Additionally, we define the similarity between the question $q$ and chunk $r_{i,j}$, calculated by embedding model $\text{g}_{i}$, as $w_{i,j}$. The differences between the Mixture-Embedding RAG and Confident RAG are detailed as follows.
3.1 Mixture-Embedding RAG
@@
In this method, we aim to directly find the chunks with the highest similarity to the question among the candidates selected by all embedding models. Intuitively, we can select chunks using the similarity score $w_{i,j}$. However, this approach can lead to two problems. First, the candidates from different embedding models may have repetitions. Therefore, we select chunks without repetition to avoid any performance loss. Second, due to the different structures of embedding models, they may have varying ranges of similarity, so that simply comparing similarity across different models may introduce potential bias. To address this, we propose to standardize the similarity score using Z-scores:
@@
3.2 Confident RAG
Inspired by Kang et al. (2025), we first let the LLM f generate NN separate answers, using the retrieved chunks from the NN embedding models separately. Then, we evaluate the confidence of the NN answers using the metrics mentioned in Section $2.$ Finally, we select the answer with the highest confidence as our final answer for each question.
4 Experiment
4.1 Experiment Setup
@@
  • Dataset: Gsm8k Cobbe et al. (2021) is a dataset of approximately 8,500 high-quality, linguistically diverse grade school math word problems designed to support question answering on basic mathematical tasks. These problems typically require 2 to 8 steps of reasoning, mainly involving elementary arithmetic operations, and are solvable by middle school students. The solutions are provided in natural language, emphasizing sequential reasoning rather than complex concepts or variables. The train dataset of Gsm8k was used as the retrieved courpus for RAG, while the first 500 items from the test dataset were used as user questions.
@@
  • Embedding models: We chose four embedding models for encoding, including: all-MiniLM-L6-v2 Face , ModernBERT-large Warner et al. (2024), MathBERT Peng et al. (2021) and stsb-roberta-large Reimers and Gurevych (2019) models.
  • LLMs: To validate our proposed method and ensure repeatability, we selected three LLMs, following Shao et al. (2025): Qwen2.5-Math-7B Yang et al. (2024), Llama-3.1-8B Grattafiori et al. (2024) and OLMo-2-1124-7B OLMo et al. (2024).
4.2 Experiment Result
4.2.1 Vanilla LLM and RAG
Table 1 presents a comparison of three different LLMs with and without RAG using different embedding models. All models achieve significant improvements in accuracy after applying RAG method, with an average improvement of 5%. The accuracy of Qwen2.5-Math-7B increases from 75.2% to 80.5%, Llama-3.1-8B from 16.6% to 21.3%, and OLMo-2-1124-7B from 21.0% to 26.0%. This demonstrates that incorporating RAG methods can effectively enhance overall model performance on the task.

### v3

# Confident RAG: Enhancing the Performance of LLMs for Mathematics Question Answering through Multi-Embedding and Confidence Scoring

@@ Abstract
Large Language Models (LLMs) hold significant promise for mathematics education, yet they often struggle with complex mathematical reasoning. While Retrieval-Augmented Generation (RAG) mitigates these issues by grounding LLMs in external knowledge, its effectiveness remains unstable, heavily dependent on the choice of a single embedding model. Moving beyond static RAG workflows, we draw on agentic workflow patterns, a paradigm that introduces structured task decomposition and collaboration to enhance system performance. We propose and examine two novel approaches that combine the benefits of multiple embedding models. While our Mixture-Embedding RAG approach (fusing retrieved documents) shows limited gains, our Confident RAG method (generating multiple answers and selecting the one with the highest confidence score) demonstrates significant improvement. Experimental results show that Confident RAG achieved average accuracy improvements of approximately 10% over vanilla LLMs and 5% over vanilla RAG. The consistent results across different LLMs and embedding models indicate that Confident RAG is an efficient plug-and-play solution for trustworthy mathematical AI assistants. Finally, we discuss how this work lays the groundwork for deploying Agentic RAG systems in educational settings, where autonomous planning and iterative refinement can be built upon our robust retrieval foundation.
I Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across various domains [lyu2025llm, zhao2025automatic, gao2024fine], showing particular promise for educational applications. However, their tendency to hallucinate [henkel2024retrieval] remains a significant barrier to reliable use in learning environments, especially in mathematics education where accuracy is crucial [frieder2023mathematical]. In mathematical reasoning, LLMs struggle with precision, often failing to process complex symbols and formulas correctly [lewkowycz2022solving], which risks misleading students when these models are deployed as educational tools.
@@ I Introduction
Retrieval-Augmented Generation (RAG) has emerged as a powerful, plug-and-play solution to mitigate hallucination by grounding LLMs in external knowledge, which has been widely employed in constructing foundation models [chen2024overview] and practical agents [arslan2024sustainable]. Unlike fine-tuning, RAG enhances model performance without extensive retraining, making it both efficient and cost-effective. The main paradigm of RAG involves first calculating the similarities between a question and chunks in an external knowledge corpus, followed by incorporating the top $K$ relevant chunks into the prompt to guide the LLMs [lewis2020retrieval].
Despite the advantages of RAG, selecting the appropriate embedding models remains a crucial concern, as the quality of retrieved references directly influences the generation results of the LLM [tu2025rbft]. Due to variations in training data and architecture, different embedding models capture semantic similarity in divergent ways. A model that excels on one type of mathematical problem may underperform on another, making the selection of a single optimal embedding model a difficult and uncertain choice. Consequently, this fundamental limitation makes vanilla RAG unreliable for comprehensive mathematics education applications.
Recent advancements in agentic workflow patterns offer new perspectives on enhancing RAG systems through structured task decomposition and collaborative processes [singh2025agentic]. These patterns provide systematic methodologies to manage complex reasoning tasks, moving beyond static retrieval pipelines toward more adaptive and robust frameworks.
@@
Inspired by these developments, we address the embedding selection challenge by proposing two methods that combine the benefits of multiple embedding models for improving RAG in mathematics education. The first method is named Mixture-Embedding RAG, which sorts the retrieved materials from multiple embedding models based on normalized similarity and selects the top $K$ materials as final references. The second method is named Confident RAG, where we first utilize vanilla RAG to generate answers multiple times, each time employing a different embedding model and recording the associated confidence metrics, and then select the answer with the highest confidence level as the final response. By validating our approach using multiple LLMs and embedding models, we illustrate the superior performance and generalization of Confident RAG, providing a reliable foundation for AI-powered mathematics education. The main contributions of this paper can be summarized as follows:
@@
  • We first point out that in RAG, different embedding models operate within their own prior domains, which is a critical barrier to deploying LLMs in mathematics education. To leverage the strengths of various embedding models, we propose and test two novel RAG methods: Mixture-Embedding RAG and Confident RAG. These methods effectively utilize the retrieved results from different embedding models to their fullest extent.
@@
II Preliminary and Related Work
II-A Retrieval Augmented Generation (RAG)
@@
Previous studies have validated a wide range of modifications to the RAG pipeline, demonstrating effectiveness in open-domain question answering by improving various components such as adaptive retrieval strategies [tang2024mba, guan2024external], dynamic document selection [kim2024re, hei2024dr], joint retriever-generator optimization [siriwardhana2021fine], reinforcement learning [huang2025rag], and hybrid retrieval-reranking frameworks [sawarkar2024blended, zhang2025levelrag].
@@
Since embedding models can produce different similarity calculations, many studies have innovated in retriever embedding models to improve the performance of the RAG pipeline. For example, Invar-RAG [liu2024invar] constructs an LLM-based retriever with LoRA-based representation learning and an invariance loss, directly addressing retriever embedding variance and locality, and demonstrates improved retrieval and response accuracy. Blended RAG [sawarkar2024blended] combines dense semantic embeddings with sparse retrievers, empirically showing that this hybrid embedding space achieves higher QA accuracy, especially as corpus size scales. W-RAG [nian2024w] further validates that enhancing retriever embeddings using weak signals from the downstream QA task offers gains comparable to methods requiring full human annotations.
However, most studies have focused on the innovation of a single embedding model. This presents a critical limitation for mathematical reasoning, where effective retrieval requires a nuanced understanding of logical structures, symbols, and problem-solving strategies. Since different embedding models capture these mathematical semantics in varied and non-uniform ways, selecting a single optimal model is inherently uncertain. Our approach will leverage the strengths of various embedding models to provide the most relevant information for LLMs.
II-B Agentic Architectures and Workflow Patterns
The limitations of static RAG workflows have catalyzed a paradigm shift towards embedding agentic intelligence into the retrieval-generation pipeline. This new paradigm, broadly categorized as Agentic RAG [ravuru2024agentic], moves beyond predefined sequences to systems capable of dynamic retrieval strategies, contextual understanding, and iterative refinement [huang2023towards].
Agentic RAG is embodied by complex architectures that integrate one or more autonomous AI agents. These architectures are built upon the foundational pillars of AI agents, which include planning for decomposing complex tasks, tool use for accessing external data and APIs, reflection for iterative self-critique and refinement, and multi-agent collaboration for specialized task handling [singh2025agentic]. Regarding current agentic RAG system, single-agent agentic RAG such as Self-RAG [asai2024self], Iter-RetGen [shao2023enhancing], and MetaRAG [zhou2024metacognitive] embed decisions about when to query external knowledge, how to structure intermediate reasoning, and when to self-critique directly into a single LLM, enabling adaptive retrieval and reflection within one policy. Plan- and graph-centric frameworks like PAR-RAG [zhang2025credible] and GraphRAG-R1 [yu2025graphrag] further externalize this agency into explicit multi-step plans or reasoning graphs that guide tool calls and evidence aggregation for multi-hop questions. In contrast, multi-agent systems such as MA-RAG [nguyen2025ma], MAO-ARAG [chen2025mao] and SIRAG [wang2025sirag] decompose the pipeline into specialized planner, retriever, evidence-selector, and answerer agents that communicate via shared memories or chain-of-thought messages, allowing more modular and interpretable control over retrieval and generation.
Agentic workflow patterns are used to support agents to optimize performance, accuracy, and efficiency under varying task complexity and resource budgets, containing prompt chaining, routing, parallelization, orchestrator-workers, and evaluator-optimizer [Anthropic2024Agents]. Prompt chaining decomposes problems into sequential subproblems, while carefully managing error propagation across steps [barrak2025traceability]. Routing directs inputs or subtasks to specialized models or agents to achieve better accuracy and cost trade-offs than any single static pipeline [zhang2025multi]. Parallelization stabilizes processing by running independent workflows concurrently, yielding more reliable solutions on complex tasks [zhang2025optimizing]. Orchestrator–worker architectures use a controller LLM to dynamically decompose tasks, delegate them to specialized executor LLMs, and coordinate their interaction, improving reliability and adaptability compared to flat single-agent setups [shi2025aime]. Finally, evaluator-optimizer iteratively critique and refine intermediate outputs, enabling smaller models to achieve larger performance at lower cost [wang2025scoreflow].
II-C Confidence of LLMs
@@
The confidence level of the LLM output can indicate the response quality to some extent. Several approaches have been employed to measure model confidence. Some of them depend on model prompting, external similarity or model-expressed judgments. These include linguistic confidence measures [shrivastava2023llamas, xiong2023can], where the model is prompted to state its confidence explicitly as part of the output, as well as LLM-as-judge or self-evaluation strategies that involve prompting the model to rate or compare its own answers in scalar or relative terms [shrivastava2025language, ren2023self, hager2025uncertainty]. However, these methods are often poorly calibrated, not directly related to the true predictive uncertainty of the model, and can be biased or inconsistent due to their reliance on language generation rather than statistical likelihood.
@@
In contrast, probability-based metrics directly ground into models’ token probability output [chen2023quantifying, jiang2021can, lin2024contextualized, chen2025first]. These include average log-probability as the mean token log-probabilities, entropy-based measures that quantify prediction uncertainty via sequence or per-token entropy, self-certainty [kang2025scalable] that utilizes the intrinsic probability distribution of LLM outputs to assess the model confidence.
@@
Many studies have observed that the probability of each token reflects the LLMs’ confidence. Intuitively, if the selected token has a higher chosen probability $p(y_{i}|x,y_{<i})$, or if the generated probability vector $\text{f}(x,y_{<i})$ exhibits a less uniform distribution, the LLMs demonstrate higher confidence in their generated results. Our study will use the confidence measurement metrics summarized in [kang2025scalable]:
@@
However, the application of these confidence metrics to mitigate retrieval-based uncertainty remains largely unexplored, particularly in specialized domains like mathematics. Here, a single error can invalidate an entire solution, making reliable confidence scoring essential. Our approach uses confidence as a criterion to select the best output from multiple embedding-grounded generations, directly addressing a key weakness of standard RAG in mathematical contexts.
II-D Research Gap and Research Questions
Previous studies reveal a clear and unaddressed opportunity. While both RAG and confidence estimation are well-studied independently, no prior work has systematically investigated the combination of multiple embedding models with confidence-based answer selection to create a robust, plug-and-play framework for mathematics. To address this gap, two research questions are proposed to guide our study:
Research question 1: Does vanilla RAG provide significant accuracy improvements over vanilla LLMs for mathematical problem-solving?
Research question 2: To what extent do the retrieval-fusion method (Mixture-Embedding RAG) and confidence-selection method (Confident RAG) enhance accuracy rates compared to standard educational AI approaches (vanilla LLM and vanilla RAG) in mathematical problem-solving?
For the Confident RAG method, we have two sub-questions:
Research question 2.1: Which confidence metrics are the most effective for selecting the best final answer in the Confident RAG framework?
Research question 2.2: What is the optimal number of embedding models for the Confident RAG framework?
III Methodology
This study addresses the challenge of unreliable AI in mathematics education by enhancing the Retrieval-Augmented Generation (RAG) framework. Based on the understanding that different embedding models may retrieve different information across the diverse spectrum of mathematical topics encountered by students, we propose two comprehensive RAG methods that combine the strengths of multiple embedding models. Both methods aim to provide students with more accurate and reliable answers by ensuring the AI is grounded in the most relevant mathematical knowledge.
III-A Workflow Overview
@@
In this work, given a student’s mathematical question $q$, we utilize a single LLM $\text{f}()$ for answer generation, while $N$ embedding models $\{\text{g}_{1}(),\text{g}_{2}(),\ldots,\text{g}_{N}()\}$ assist in retrieving question-related information from a corpus involving mathematics textbooks and question-answer pairs, which has been divided into chunks $C=[c_{1},c_{2},\ldots,c_{m}]$. This setup mirrors a scenario where a student seeks help from a tutor who consults multiple specialized reference books. Inspired by the stacking and bagging paradigms of ensemble learning, we propose two distinct methods: (1) Mixture-Embedding RAG, which first combines the retrieved results from multiple embedding models and incorporates them into the prompt for the LLM, and (2) Confident RAG, which employs vanilla RAG with different embedding models for initial answer generation and subsequently selects the answer with the highest confidence as the final response.
@@
As shown in Fig. 1, the two proposed methods share the same preliminary process: first, they utilize different embedding models to identify the $k$ chunks with the highest similarity, similar to the approach taken by the vanilla RAG described in Section II-A. We define the retrieved chunks as $R=\{r_{i,j}\mid i=1,2,\ldots,N;j=1,2,\ldots,k\}$, where $r_{i,j}$ is defined as $c_{\mathcal{K}_{j}^{i}}$ and $\mathcal{K}^{i}$ is the index set of the selected $k$ chunks by embedding model $\text{g}_{i}$. Additionally, we define the similarity between the question $q$ and chunk $r_{i,j}$, calculated by embedding model $\text{g}_{i}$, as $w_{i,j}$. The differences between the Mixture-Embedding RAG and Confident RAG are detailed as follows.
III-B Mixture-Embedding RAG
@@
In this method, we aim to directly find the chunks with the highest similarity to the question among the candidates selected by all embedding models. The goal is to mimic an expert tutor who synthesizes information from multiple sources to explain a concept. Intuitively, we can select chunks using the similarity score $w_{i,j}$. However, this approach can lead to two problems. First, the candidates from different embedding models may have repetitions. Therefore, we select chunks without repetition to avoid any performance loss. Second, due to the different structures of embedding models, they may have varying ranges of similarity, so that simply comparing similarity across different models may introduce potential bias. To address this, we propose to standardize the similarity score using Z-scores:
@@
III-C Confident RAG
Inspired by [kang2025scalable], we first let the LLM f generate $N$ separate answers, using the retrieved chunks from the $N$ embedding models separately. Then, we evaluate the confidence of the $N$ answers using the metrics mentioned in Section II-C. Finally, we select the answer with the highest confidence as our final answer for each question. This method is particularly valuable in educational contexts where a single incorrect solution can significantly impact student learning. By selecting the answer the LLM is most certain about, we provide a safer, more reliable mathematical assistant for students.
IV Experiment
IV-A Experiment Setup
@@
  • Dataset: Gsm8k [cobbe2021gsm8k] is a dataset of approximately 8,500 high-quality, linguistically diverse grade school math word problems designed to support question answering on basic mathematical tasks. These problems typically require 2 to 8 steps of reasoning, mainly involving elementary arithmetic operations, and are solvable by middle school students. The solutions are provided in natural language, emphasizing sequential reasoning rather than complex concepts or variables. The train dataset of Gsm8k was used as the retrieved courpus for RAG, while the first 500 items from the test dataset were used as user questions.
@@
  • Embedding models: We chose four embedding models for encoding, including: all-MiniLM-L6-v2 [all-mini], ModernBERT-large [modernbert], MathBERT [peng2021mathbert] and stsb-roberta-large [reimers-2019-sentence-bert] models.
  • LLMs: To validate our proposed method and ensure repeatability, we selected three LLMs, following [shao2025spuriousrewardsrethinkingtraining]: Qwen2.5-Math-7B [yang2024qwen2], Llama-3.1-8B [grattafiori2024llama] and OLMo-2-1124-7B [olmo20242].
IV-B Experiment Result
IV-B1 Does vanilla RAG provide significant accuracy improvements over vanilla LLMs for mathematical problem-solving?
Table I presents a comparison of three different LLMs with and without RAG using different embedding models. All models achieve significant improvements in accuracy after applying RAG method, with an average improvement of 5%. The accuracy of Qwen2.5-Math-7B increases from 75.2% to 80.5%, Llama-3.1-8B from 16.6% to 21.3%, and OLMo-2-1124-7B from 21.0% to 26.0%. This confirms that grounding LLMs in external knowledge is a viable strategy for improving their reliability in mathematical contexts.
However, the performance of vanilla RAG was unstable when using different embedding models. This limitation can be vividly illustrated in Table II for a 400-meter hurdle word problem. The vanilla LLM, without any retrieved context, produced an incorrect answer of 34.2. When augmented with RAG, the results were different: using embedding model 1 or 4 led to the same incorrect answer (34.2), while models 2 and 3 successfully produce the right answer (36).
TABLE I: Model Performance Comparison Across Different Embedding Methods
LLM Model Vanilla LLM Vanilla RAG
Emb1 Emb2 Emb3 Emb4 Avg Improvement